OpenStudy (anonymous):

The mass of a sphere M varies directly as the cube of the radius r. If M = 0.54 kilograms when r = 3 centimeters, what is the mass of the sphere, made of the same material, with a radius of 5 centimeters?

OpenStudy (whpalmer4):

\[M = kr^3\]If you increase the radius by a factor of 5/3, the mass changes by a factor of (5/3)^3.

OpenStudy (anonymous):

2.5 kilograms is what i got but not entirely sure

OpenStudy (whpalmer4):

You should be sure :-)

OpenStudy (anonymous):

Thanks

OpenStudy (whpalmer4):

This may not be the way they expected you to solve this (probably wanted you to solve \(M = kr^3\) for \(k\) when you know \(M\) and \(r\) and use that to find the new M\) but for these scaling problems, just replace the dimensions with the scaling factors to get the overall change. Surface area of a sphere when the radius doubles? Well, surface area of sphere is proportional to r^2, so (2)^2 = 4, surface area doubles. Etc.

OpenStudy (whpalmer4):

Surface area quadruples, that is, not doubles...

OpenStudy (whpalmer4):

Or more usefully, radius of pizza doubles, therefore area of pizza goes up by (2)^2 = 4 :-)

OpenStudy (anonymous):

Thanks alot for that information